# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ 8 Implementation
We implemented all experiments, including the training of SBN and downstream applications, using the PyTorch framework paszke2019pytorch. The SBN model was trained in a supervised manner, with numerical solutions from the Least Squares Quasiconformal Energy (LSQC) solver serving as ground-truth mappings. The training dataset comprised 960 images from the ILSVRC2012 dataset russakovsky2015imagenet, which provided realistic BCs with both low- and high-frequency components. To generate BCs, we randomly selected two images per training iteration, applied data augmentation techniques (random flipping, noise addition, and blurring), and normalized them to form the real and imaginary parts of the BCs. Two pinned points were randomly chosen from the vertices of the finest mesh for each training sample.
SBN employs a hierarchical architecture. It integrates two core components: multiscale message-passing mechanism and mesh spectral layer, as detailed in Sec.6. The numbers of message passing numbers per scale depends on the mesh resolution, which are 3,3,5 from high to low, and the number of MSL per scale is 2. The training process ran for 250 epochs using the AdamW optimizer loshchilov2017decoupled with parameters $\beta=(0.9,0.999)$, a weight decay of $1\times 10^{-2}$, an inital learning rate $3\times 10^{-4}$. We used three triangular meshes of varying resolutions, generated with the MATLAB DistMesh function persson2004simple, with initial edge lengths of 0.015, 0.3, and 0.5. The numbers of vertices and faces in these three meshes with resolution from high to low are (16114, 31803), (4027,7846) and (1452,2774), respectively. The loss function for training SBN combined three terms $\mathcal{L}_{i}^{ds}(i=1,2,3)$ with weights all 1. Although we did not perform an exhaustive grid search for optimal hyperparameters, this configuration yielded strong performance, as demonstrated in Sec.9. All experiments were conducted on an NVIDIA A40 GPU with 44GB memory, ensuring efficient computation.
@@ 9 Experimental results
Numerical experiments were conducted to evaluate the efficacy and potential of SBN. Given that SBN’s performance depends on its ability to accurately predict mappings for various BCs $\mu$, we first compare its predicted outputs with numerical solutions. This is followed by several applications in density equalizng and inconsistent surface registration, encompassing both realistic and synthetic scenarios.
@@ 9.1 Performance of Spectral Beltrami Network
SBN, trained over 250 epochs, achieves robust simulation of the Beltrami equation. Under typical conditions (BC with norms mostly below 0.4), it achieves the average nodewise L2 error smaller than 0.03, relative to numerical LSQC solutions, demonstrating strong generalizability on unseen data (Fig.5). The model maintains robustness in challenging scenarios involving larger BC norms (Fig.6) or significant domain distortions (Fig.7), effectively handling complex or large deformations.
Figure 5: Comparison of SBN, M-MGN, and MSN for typical cases (max BC norm ¡ 0.4). Top row: numerical solution field (left), BC magnitude (center), log gradient magnitude of BCs (right). Bottom row: color-coded nodewise L2 errors for SBN, M-MGN, and MSN vs. reference solution. Format applies to Figs.6 and 7.
@@
Figure 6: Comparison of three models (SBN, M-MGN and MSN) for simulating LSQC in the case that BCs’ range is larger.
Figure 7: Comparison of three models (SBN, M-MGN and MSN) for simulating LSQC in the case that the desired mapping is a large distortion.
@@
A self-ablation study was conducted to investigate individual contributions of the multiscale message-passing mechanism and the mesh spectral layer (MSL) within our architecture. Removing MSL yields Multiscale MeshGraphNet (M-MGN) fortunato2022multiscale; using only MSL gives Mesh Spectral Net (MSN). Across all three test cases (Figs.5-7), SBN significantly outperforms both ablations, confirming the synergy between multiscale message passing and spectral layers.
@@
To further evaluate SBN, we compared it with M-MGN, MSN, and the numerical LSQC algorithm, running tests across 10,000 sets of BCs and pinned points. Two regimes were studied: (1) maximum BC L2 norm less than 0.4, and (2) maximum BC L2 norm between 0.4 and 0.8. We measure: non-injective triangles (faces with negative Jacobian), nodewise L2 error, boundary L2 error, and BC L2 error (vs. target face-averaged BCs). Tables 12 show that SBN achieves the lowest errors and fewest fold-overs, closely matching numerical LSQC accuracy while substantially outperforming M-MGN and MSN in both regimes.
Table 1: Performance for Case 1 (maximum BC norm is smaller than 0.4), averaged over 10,000 tests. Lower values indicate better performance.
Method Non-injective Triangles Nodewise L2 Error Boundary L2 Error BC L2 Error
LSQC 0.00050 0.00348
SBN 0.09560 0.02239 0.03352 0.02306
M-MGN 0.38050 0.09481 0.15194 0.07031
MSN 33.2049 0.49009 0.73313 0.09514
@@
Table 2: Performance for Case 2 (max BC norm is between 0.4 and 0.8), averaged over 10,000 tests. Lower values indicate better performance.
Method Non-injective Triangles Nodewise L2 Error Boundary L2 Error BC L2 Error
LSQC 0.24530 0.01146
SBN 1.28710 0.04671 0.07217 0.07679
M-MGN 23.5595 0.23172 0.37515 0.19477
MSN 264.365 0.47382 0.70073 0.18091
9.2 Density Equalizing parameterization
We demonstrate SBN-Opt’s superiority over numerical methodschoi2018density in both density equalization and conformal distortion control. Problem formulation details are in Sec.7.
@@ 9.2.1 Simply Connected Cases
We begin with two simply connected examples. First is a synthetic example of a peak. In this example, the weights of task loss $E_{1}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 5e-2 and 1e-3, respectively. The result mappings of our method and DEMchoi2018density are shown in Fig.8. In Fig.9, we compared our result with DEM’s result in two aspects, distribution of angle distortion and performance in density equalizing $\rho(T)=\frac{\text{Area}(f(T))}{\text{Area}(T)}$. Our SBN-Opt achieved a more uniform density distribution, reducing the variance of face density from 0.7382 to a mere 0.0017. Simultaneously, it produced a mapping with lower angular distortion, as evidenced by a 3.5% reduction in the average L2 norm of the resulting BCs (from 0.481 to 0.464).
Figure 8: Density equalizing parameterization of the ”peak” model. From left to right: The original 3D mesh, the parameterization result from our SBN-Opt, and the result from the DEM algorithm. All meshes are colored according to face area, where warmer colors indicate larger areas.
@@
Figure 9: Statistical comparison of parameterization results for the ”peak” model. From left to right: (a) Histogram of BC magnitudes from our method; (b) Histogram of BC magnitudes from DEM; (c) Histogram of the face density distribution from our method; and (d) Histogram of the face density distribution from DEM. Note that the histogram for the DEM density distribution (d) is truncated to the density range of our method’s result to enable a direct comparison, as the DEM result contains significant outliers.
The second example is a lion face. In this example, the weights of task loss $E_{1}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 1e-2 and 1e-3, respectively. The result mappings of our method and DEM are shown in Fig.10. As shown in Fig.9, the improvements were even more pronounced. Our method decreased the average L2 norm of the BCs by 11.6% (from 0.405 to 0.358), indicating substantially less conformal distortion than the DEM result. The final density distribution was also highly concentrated around the target, with its variance decreasing significantly to 0.0062.
Figure 10: Density equalizing parameterization of the ”lion face” model. From left to right: The original 3D mesh, the parameterization result from our SBN-Opt, and the result from the DEM algorithm. All meshes are colored according to mean curvature for visualization.
Figure 11: Quantitative comparison for the ”lion face” model against the DEM algorithm. The meaning of each histogram corresponds to that in Fig.9.
Across all cases, SBN-Opt consistently outperforms DEM in both density equalization and angular distortion reduction. DEM’s higher variance stems from outlier faces with extreme densities; even excluding these, DEM’s variance exceeds ours, confirming SBN-Opt’s superior global uniformity.
@@ 9.2.2 Multiply Connected Cases
Our method extends seamlessly to multiply connected domains. The first example is a human face with one hole. In this example, the weights of task loss $E_{1}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 5e-1 and 1e-4, respectively. Let $S$ be a surface, $f:S\rightarrow\mathbb{C}$ be the density-equalizing map, and considering a set of uniformly distributed points $\mathcal{P}$ on $f(S)$, we can triangulate them and obtain a new triangulation $\mathcal{T}$. Then, using the inverse mapping $f^{-1}$, we can interpolate $\mathcal{P}$ onto $S$ and obtain a remeshed triangulation $(f^{-1}(\mathcal{P}),\mathcal{T})$ of the surface $S$. The results are shown in Fig.12. The average L2 norm of BCs induced by the result mapping is 0.097 and the variance of density is 0.0084.
The second example is a human face with two holes. In this example, the weights of task loss $E_{1}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 5e-1 and 1e-4, respectively. The results are shown in Fig.12. The average L2 norm of BCs induced by the result mapping is 0.124 and the variance of density is 0.0127.
[ Density equalizing parameterization for the human face with one hole.] Refer to caption  [Density equalizing parameterization for the human face with two holes.] Refer to caption

### v2

@@ 7 Implementation
We implemented all experiments, including the training of SBN and downstream applications, using the PyTorch framework Paszke (2019). The SBN model was trained in a supervised manner, with numerical solutions from the Least-squares Quasiconformal Energy (LSQC) solver serving as ground-truth mappings. The training dataset comprised 960 images from the ILSVRC2012 dataset Russakovsky et al. (2015), which provided realistic BCs with both low- and high-frequency components. To generate BCs, we randomly selected two images per training iteration, applied data augmentation techniques (random flipping, noise addition, and blurring), and normalized them to form the real and imaginary parts of the BCs. Two pinned points were randomly chosen from the vertices of the finest mesh for each training sample.
SBN employs a hierarchical architecture. It integrates two core components: multiscale message-passing mechanism and mesh spectral layer, as detailed in Section 5. The numbers of message passing numbers per scale depends on the mesh resolution, which are 3,3,5 from high to low, and the number of MSL per scale is 2. The training process ran for 250 epochs using the AdamW optimizer Loshchilov and Hutter (2017) with parameters $\beta=(0.9,0.999)$, a weight decay of $1\times 10^{-2}$, an inital learning rate $3\times 10^{-4}$. We used three triangular meshes of varying resolutions, generated with the MATLAB DistMesh function Persson and Strang (2004), with initial edge lengths of 0.015, 0.3, and 0.5. The numbers of vertices and faces in these three meshes with resolution from high to low are (16114, 31803), (4027,7846) and (1452,2774), respectively. The loss function for training SBN combined three terms $\mathcal{L}_{i}^{ds}(i=1,2,3)$ with weights all 1. Although we did not perform an exhaustive grid search for optimal hyperparameters, this configuration yielded strong performance, as demonstrated in Section 8. All experiments were conducted on an NVIDIA A40 GPU with 44GB memory, ensuring efficient computation. In addition, the numerical algorithms were implemented using MATLAB on an 8-core Intel i7 machine.
@@ 8 Experimental results
Numerical experiments were conducted to evaluate the efficacy and potential of SBN. Given that SBN’s performance depends on its ability to accurately predict mappings for various BCs $\mu$, we first compare its predicted outputs with numerical solutions. This is followed by several applications in equiareal parameterization and inconsistent surface registration, encompassing both realistic and synthetic scenarios.
@@ 8.1 Performance of Spectral Beltrami Network
SBN, trained over 250 epochs, achieves robust simulation of the Beltrami equation. Under typical conditions (BC with norms mostly below 0.4), it achieves the average nodewise $\ell^{2}$ error smaller than 0.03, relative to numerical LSQC solutions, demonstrating strong generalizability on unseen data (Figure 4). The model maintains robustness in challenging scenarios involving larger BC norms (Figure 5) or significant domain distortions (Figure 6), effectively handling complex or large deformations.
Figure 4: Comparison of SBN, M-MGN, and MSN for simulating LSQC in a typical case where the maximum Beltrami coefficient norm is less than 0.4. Top row: ground truth mapping (left) and the magnitude of the BC of the ground truth mapping (right). Bottom row: results of SBN, M-MGN, and MSN, respectively. For each model, we report the average nodewise $\ell^{2}$ error with respect to the numerical LSQC solution and the number of flipped triangles (faces with negative Jacobian) above the corresponding subfigure. Same format applies to Figure 5 and Figure 6.
@@
Figure 5: Comparison of three models (SBN, M-MGN and MSN) for simulating LSQC when the prescribed Beltrami coefficients have a larger range.
Figure 6: Comparison of SBN, M-MGN, and MSN for simulating LSQC in a case where the desired mapping exhibits a large distortion. In addition, four red rectangular frames are drawn to highlight regions of interest: the red frames in the M-MGN and MSN results enclose areas containing flipped triangles, visualized as light red triangles in the zoomed-in views; the corresponding red frames in the SBN result mark the same regions on our mapping, where the triangles remain non-flipped and well-shaped.
@@
A self-ablation study was conducted to investigate individual contributions of the multiscale message-passing mechanism and the mesh spectral layer (MSL) within our architecture. Removing MSL yields Multiscale MeshGraphNet (M-MGN) Fortunato et al. (2022); using only MSL gives Mesh Spectral Net (MSN). Across all three test cases (Figure 4-Figure 6), SBN significantly outperforms both ablations, confirming the synergy between multiscale message passing and spectral layers.
@@
To further evaluate SBN, we compared it with M-MGN, MSN, and the numerical LSQC algorithm, running tests across 10,000 sets of BCs and pinned points. Two regimes were studied: (1) maximum BC $\ell^{2}$ norm less than 0.4, and (2) maximum BC $\ell^{2}$ norm between 0.4 and 1. We measured: non-injective triangles (faces with negative Jacobian), nodewise $\ell^{2}$ error, boundary $\ell^{2}$ error, and BC $\ell^{2}$ error (vs. target face-averaged BCs). Table 1(a) and Table 1(b) show that SBN achieves the lowest approximation errors among 3 models in both regimes, compared with numerical results.
(a) Case 1 :maximum BC norm is smaller than 0.4
Method Nodewise $\ell^{2}$ Error Boundary $\ell^{2}$ Error BC $\ell^{2}$ Error
SBN 0.02239 0.03352 0.02306
M-MGN 0.09481 0.15194 0.07031
MSN 0.49009 0.73313 0.09514
(b) Case 2: max BC norm is between 0.4 and 1
Method Nodewise $\ell^{2}$ Error Boundary $\ell^{2}$ Error BC $\ell^{2}$ Error
SBN 0.04671 0.07217 0.07679
M-MGN 0.23172 0.37515 0.19477
MSN 0.47382 0.70073 0.18091
@@
Table 1: Performance of SBN, M-MGN and MSN in two cases, both averaged over 10,000 tests. Lower values indicate better performance.
8.2 Equiareal parameterization
We demonstrated SBN-Opt’s superiority over numerical methods DEMChoi and Rycroft (2018) and DEQ Lyu et al. (2024a) in both area preserving and conformal distortion control. Problem formulation details are in Section 6. Our assessment was performed in both simply and multiply connected scenarios, systematically examining three aspects: the distribution of BCs, face density uniformity, and angle distortion of 2D parameterization result compared with 3D surface. The angle distortion of a triangle $T_{i}$ in the 2D parameterization is defined as $A_{i}=\frac{1}{3}\sum_{j=1}^{3}|\theta_{ij}-\tilde{\theta}_{ij}|$ where $\theta_{ij},\tilde{\theta}_{ij}$ are $j$-th angle of $T_{i}$ in the parameterization and its correspondence $U_{i}$ in the 3D surface, respectively.
@@ 8.2.1 Simply Connected Cases
We first considered two simply connected examples: a synthetic “peak” model and a lion head surface. In both experiments, we compared our SBN-Opt framework with the DEM method Choi and Rycroft (2018).
For the peak model, we set the weights of the task loss $E_{1}$, the angle distortion regularizer $E_{\text{BC}}$, and the smoothness regularizer $E_{\text{smooth}}$ to 1, 5e-2 and 1e-3, respectively. The resulting parameterizations are shown in Figure 7 with each face colored according to the density error $\frac{|\text{area}(T)-\text{area}(f(T))|}{|\text{area}(f(T))|}$. Indicated by Figure 8, our SBN-Opt produced an equiareal parameterization whose face-area distribution is highly concentrated around the target density, with the variance of the face density reduced from 0.7382 to 0.0052. At the same time, SBN-Opt achieved lower conformal distortion: the mean BC norm decreased from 0.2455 to 0.2418, and the mean angle distortion decreased from $29.79^{\circ}$ to $28.57^{\circ}$. Importantly, our mapping remained a homeomorphism without any overlaps while DEM’s result contained 4 flipped triangles. The histograms in Figure 8 illustrate that, compared with DEM, SBN-Opt notably suppressed extreme outliers in the density distribution while improving angular fidelity.
Figure 7: Equiareal parameterization of the ”peak” model. From left to right: The original 3D mesh, the parameterization result from our SBN-Opt, and the result from the DEM algorithm. The 2D parameterization results are colored associated with the relative error $\frac{|\text{area}(T)-\text{area}(f(T))|}{|\text{area}(f(T))|}$, and the overall color of SBN-Opt’s result is deeper than that of DEM’s result, which means SBN-Opt’s result is more area-preserving.
@@
Figure 8: Statistical comparison of parameterization results for the peak model. Top row: results for DEM; bottom row: results for SBN-Opt. From left to right: (a) Histogram of BC magnitudes; (b) Histogram of per-face angle distortion (in degrees); (c) Histogram of the face density distribution. Note that the density histogram of DEM’s result is visibly truncated to the value range of SBN-Opt’s result so that the two distributions can be compared on the same scale. Same format applies to Figure 10, Figure 13 and Figure 14.
In the second example, the weights of task loss $E_{1}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ were 1, 1e-2 and 1e-2, respectively. Figure 9 shows the parameterization results, and Figure 10 reports the corresponding statistics. Compared with DEM, SBN-Opt substantially reduced conformal distortion: the average BC norm was reduced by 26.0% (from 0.4599 to 0.3402), and the mean angle distortion decreased by 10.5% (from $23.24^{\circ}$ to $20.81^{\circ}$). The area preserving performance was also greatly improved, with the variance of the face density reduced to 0.0074, whereas there were faces with significantly distorted densities in the DEM’s result. Again, no overlaps were observed in our result. Overall, on simply connected surfaces, SBN-Opt consistently produced maps that are both closer to the target density and more conformal than DEM.
Figure 9: Left: original lion head mesh shown from front (top) and back (bottom). Middle: DEM’s density-equalizing parameterization. Two red dashed rectangles mark regions are zoomed in where the remeshed triangles are severely distorted and contain local fold-overs (flipping triangles). Right: SBN-Opt’s equiareal parameterization. The red dashed rectangles indicate the regions corresponding to those highlighted in the DEM result; the zoom-in views show that the triangles in these areas are non-flipping and well shaped.
Figure 10: Statistical comparison of parameterization results for the lion head model.
Figure 11: Remeshing comparison for the lion head surface at two resolutions using DEM and SBN-Opt. From left to right: remeshed result by DEM with 4.6K vertices, remeshed result by SBN-Opt with 4.6K vertices, remeshed result by DEM with 30K vertices, and remeshed result by SBN-Opt with 30K vertices. For each method and resolution, the top row shows a front view of the remeshed surface and the bottom row shows a back view. Red rectangular boxes mark regions of interest near the lower part of the mesh, which implies that SBN-Opt produces more regular and uniformly sized elements in the highlighted regions than DEM.
@@ 8.2.2 Multiply Connected Cases
We next tested SBN-Opt on multiply connected surfaces and compared it with the DEQ methodLyu et al. (2024a). We considered two examples: a human face with one hole and a human face with two holes.
For the human face with one hole, we set the weights of $E_{1}$, $E_{\text{BC}}$ and $E_{\text{smooth}}$ to 1, 1e-2 and 1e-2, respectively. The parameterization obtained by SBN-Opt and DEQ are shown in Figure 13, with quantitative statistics in Figure 13. Our method achieved a noticeable improvement in conformal quality: the average BC norm was reduced by 15.8% (from 0.1416 to 0.1192), and the mean angle distortion decreased by 17.0% (from $8.208^{\circ}$ to $6.810^{\circ}$). At the same time, area preserving performance was significantly enhanced: the variance of the face density dropped from 7.996e-2 for DEQ to 9.47e-4 for SBN-Opt.
In the second example, the weights of task loss $E_{1}$, $E_{\text{BC}}$ and $E_{\text{smooth}}$ were 1, 1.5e-1 and 0, respectively. The visual comparison is given in Figure 12(b) and the statistical comparison between our result and DEQ’s is shown in Figure 14. Again, SBN-Opt yielded a parameterization with substantially reduced conformal distortion: the average Beltrami norm decreased by 15.5% (from 0.1305 to 0.1095), and the mean angle distortion decreased by 13.5% (from $7.244^{\circ}$ to $6.270^{\circ}$). In terms of area distribution, our method produced a density that is much closer to the target, with its variance reduced from 31.46 to 1.517e-2.
Across both multiply connected examples, SBN-Opt consistently delivered parameterizations that better balance area preserving and conformal distortion than DEQ. By directly optimizing over Beltrami coefficients, our framework avoided the severe density outliers and excessive angular distortion that appear in the numerical baselines.
@@
Figure 13: Statistical comparison of parameterization results for the human face with one hole.
Figure 14: Statistical comparison of parameterization results for the human face with two holes.
Figure 15: Remeshing comparison for human face surfaces with one (left) and two holes (right), respectively. For each example, the top row shows the remeshed result obtained from the baseline method (DEQ), and the bottom row shows the remeshed result obtained from the SBN-Opt–based parameterization. Around each face, red dashed rectangles indicate zoom-in windows placed along the eye or nose boundaries; the inset views display the local triangle patterns in these regions. The side-by-side zoomed comparisons reveal that the SBN-Opt based remeshing yields more regular, well-shaped, and uniformly sized elements.
8.2.3 Application to Remeshing
A high-quality equiareal parameterization is not only an end in itself, but also a good tool for downstream geometry processing. In this subsection, we demonstrate how the equiareal parameterization produced by SBN-Opt can be used to generate high-quality remeshed 3D surfaces, and how the improved parameterization quality directly translates into better remeshing results.
Let $S$ be a surface, $f:S\rightarrow\mathbb{C}$ be the equiareal map, and considering a set of uniformly distributed points $\mathcal{P}$ on $f(S)$, we can triangulate them and obtain a new triangulation $\mathcal{T}$. Then, using the inverse mapping $f^{-1}$, we can interpolate $\mathcal{P}$ onto $S$ and obtain a remeshed triangulation $(f^{-1}(\mathcal{P}),\mathcal{T})$ of the surface $S$.To assess how the quality of the parameterization impacts the remeshing, we applied this procedure to three surfaces in the last experiments and obtain: lion head with 4.6K vertices (see Figure 11 left), lion head with 30K vertices (see Figure 11 right), human face with one hole (see Figure 15(a)) and human face with two holes (see Figure 15(b)). For each surface, we constructed remeshed results from the parameterizations produced by SBN-Opt and by the baseline DEM/DEQ methods. In Figure 11, Figure 15(b) and Figure 15(b), we highlight several zoomed-in regions with red triangular frames. Visually, the remeshed surfaces derived from our SBN-Opt parameterizations exhibited more regular triangle shapes and more uniform sampling density, while the DEM/DEQ-based remeshed surfaces contained noticeably skewed and stretched triangles in regions where the underlying parameterizations suffered from larger distortion.
To quantify these observations, in addition to the angle distortion between 2D uniform meshes and 3D remeshed surfaces, we evaluated two more complementary metrics on the remeshed 3D surfaces:
  1. 1.
    Size variation. On the remeshed 3D surface $\widetilde{\mathcal{M}}$, we computed the standard deviation of the triangle areas:
    $\delta_{\text{size}}=\text{std}\bigl(\text{Area}(T)\bigr),$
    where $\text{Area}(T)$ is the area of each face $T$ in $\widetilde{\mathcal{M}}$. A smaller $\delta_{\text{size}}$ reflects a more uniform distribution of triangle sizes, which is consistent with a better equiareal parameterization.
  2. 2.
    Shape variation For each triangle $T_{i}$ in $\widetilde{\mathcal{M}}$, let $e_{ij}$ be the length of its three edges. We measured the deviation of the normalized edge lengths from the equilateral configuration:
    $R_{i}=\sum_{j=1}^{3}\left|\frac{e_{ij}}{\sum_{j=1}^{3}e_{ij}}-\frac{1}{3}\right|.$
    The shape variation is then defined as
    $\delta_{\text{shape}}=\text{mean}(R_{i}).$
    A smaller $\delta_{\text{shape}}$ value indicates better control of angular distortion in the remeshing.
For each of the four remeshing examples, we computed these three metrics for the remeshed surfaces obtained from SBN-Opt and from DEM/DEQ, and summarized the results in the Table 2.
Table 2: The performance of remeshed Results by SBN-Opt and DEM. We recorded the mean angle distortion, size variation and shape variation of each result.
Subject Method Mean angle distortion Size variation Shape variation
Lion Head(4.6K) DEM 21.621582 0.2355 0.6022
SBN-Opt 21.35 0.1975 0.1182
Lion Head(30K) DEM 23.132975 0.2219 0.1960
SBN-Opt 22.30 0.2023 0.01620
Human face 1 hole(10K) DEQ 7.586 0.07717 2.2819e-5
SBN-Opt 7.139 0.07566 1.398e-5
Human face 2 holes(10K) DEQ 6.924 0.07162 2.5748e-5
SBN-Opt 6.755 0.07124 2.000e-5
Across all cases, the SBN-Opt–based remeshes consistently exhibited lower average angle distortion, smaller area variation, and reduced shape variation, compared with the DEM/DEQ-based remeshes. These quantitative improvements aligned with the visual comparisons in Figure 11, Figure 15(a) and Figure 15(b): triangles generated from our parameterizations were more regular and more uniformly sized, particularly in regions of high curvature or around holes where DEM/DEQ tends to produced elongated or crowded elements.
In summary, these experiments confirmed that a parameterization with lower angular distortion and better area preserving is crucial for high-quality remeshing. By explicitly optimizing over Beltrami coefficients and directly controlling local distortion, SBN-Opt not only outperforms numerical methods at the parameterization level, but also yields superior remeshed 3D surfaces in downstream applications.